AS/AD Model.
AD: P = 48 - 2Y.
AS: P = 8.
PO: Y = 16.
What is the Inflation??
1
Expert's answer
2016-05-15T12:38:02-0400
If in AS/AD Model AD function is P = 48 - 2Y, AS function is P = 8, potential output (PO) is Y = 16, then in equilibrium:
AS = AD, so 48 - 2Y = 8, Y = 20.
If at PO Y = 30, then there is an inflationary gap (real GDP is higher than potential GDP), so there is an inflation in this case.
Inflation rate is pi = (Yr/Yp - 1)*100% = (20/16 - 1)*100% = 25%.